### v1

# Steiner trees with infinitely many terminals on the sides of an angle

Abstract
The Euclidean Steiner problem is the problem of finding a set ${\mathcal{S}t}$, with the shortest length, such that ${\mathcal{S}t}\cup\mathcal{A}$ is connected, where $\mathcal{A}$ is a given set in a Euclidean space. The solutions ${\mathcal{S}t}$ to the Steiner problem will be called Steiner sets while the set $\mathcal{A}$ will be called input. Since every Steiner set is acyclic we call it Steiner tree in the case when it is connected. We say that a Steiner tree is indecomposable if it does not contain any Steiner tree for a subset of the input.
We are interested in finding the Steiner set when the input consists of infinitely many points distributed on two lines. In particular we would like to find a configuration which gives an indecomposable Steiner tree.
It is natural to consider a self-similar input, namely the set $\mathcal{A}_{\alpha,\lambda}$ of points with coordinates $(\lambda^{k-1}\cos\alpha,$$\pm\lambda^{k-1}\sin\alpha)$, where $\lambda>0$ and $\alpha>0$ are small fixed values. These points are distributed on the two sides of an angle of size $2\alpha$ in such a way that the distances from the points to the vertex of the angle are in a geometric progression.
To our surprise, we show that in this case the solutions to the Steiner problem for $\mathcal{A}_{\alpha,\lambda}$, when $\alpha$ and $\lambda$ are small enough, are always decomposable trees. More precisely, any Steiner tree for $\mathcal{A}_{\alpha,\lambda}$ is a countable union of Steiner trees, each one connecting 5 points from the input. Each component of the decomposition can be mirrored with respect to the angle bisector providing $2^{\mathbb{N}}$ different solutions with the same length. By considering only a finite number of components we obtain many solutions to the Steiner problem for finite sets composed of $14k+$ 1 points distributed on the two lines ($12k+$ 1 on a line and $2k$ on the other line). These solutions are very similar to the ladders of Chung and Graham.
We are able to obtain an indecomposable Steiner tree by adding, to the previous input, a single point strategically placed inside the angle. In this case the solution is in fact a self-similar tree (in the sense that it contains a homothetic copy of itself).
Finally, we show how the position of the Steiner points in the Steiner tree can be described by a discrete dynamical system which turns out to be equivalent to a 2-interval piecewise linear contraction.
1 Introduction
The finite Euclidean Steiner problem is the problem of finding a one-dimensional connected set ${\mathcal{S}t}$ of minimal length containing a finite set of given points $\mathcal{A}\subset\mathbb{R}^{d}$.
The history of the finite Euclidean Steiner problem is studied in the paper [2]. Brazil, Graham, Thomas and Zachariasen did a detailed research and discovered that the statement and basic results about the Steiner Problem were rediscovered (at least) three times: it was first stated by Gergonne in 1811, by Gauss in 1836 and by Jarník and Kössler [14] in 1934. The problem has become well-known as the “Steiner problem” after the great success of the book “What is Mathematics?” by Courant and Robbins [8].
In 1980-s and 1990-s, explicit solutions to the finite Steiner problem attracted the attention of several notable mathematicians. It is worth noting that Du, Hwang and Weng [9] completely solved the Steiner problem when $\mathcal{A}$ is the set of vertices of a regular polygon. Rubinstein and Thomas [24] generalizes the result when the points of $\mathcal{A}$ are uniformly enough distributed along a circle.
A setting which is similar to ours, can be found in the paper by Burkard and Dudás [3], who determined the Steiner trees for the vertex of a plane angle of size at least $\pi/6$ and all points in the sides at a distance $1,\dots,n$. It turns out that apart from the points which are close to the vertex, the Steiner tree is composed of segments of length $1$ joining the points on the sides of the angle. The results of our paper are also very similar to the results about ladders of Chung and Graham [7] and Burkard, Dudás and Maier [4]. A ladder is a collection of $2n$ lattice points placed on the vertices of $n-1$ adjacent unit squares, forming a rectangle of sides $1\times(n-1)$. It turns out that when $n$ is odd the Steiner tree is full, which means that terminal points are reached by a single edge of the tree. While if $n$ is odd the Steiner tree is not full. By splitting the tree in the terminals of order two one can decompose the Steiner tree into the union of small Steiner trees, each one containing only two or four terminal points. Notice that this is a case (relevant to our quest) where a full Steiner tree with an arbitrarily large number of points is presented. However, as the number of points is going to infinity also the length of the Steiner tree is going to infinity, hence this example is not suitable to construct a tree with infinitely many vertices.
Finding the Steiner tree for a finite set of points is a well-known problem in computational geometry. Garey, Graham and Johnson [11] proved that the Steiner problem is NP-hard. Rubinstein, Thomas and Wormald [25] proved that such a complexity persists even when the terminals are constrained to lay along two parallel lines and also in the case when the terminals are placed on the sides of an angle with size which is smaller than $32\pi/$ 3 (see Theorem 3 in [25]).
More recently, a generalized setting for the Steiner problem was given by Paolini and Stepanov [19]: the ambient space $X$ can be any connected complete metric space with the Heine–Borel property (closed bounded sets are compact) and the given set of points can be any compact subset of the ambient space. In this setting there always exists a set ${\mathcal{S}t}$, with minimal $1$-dimensional Hausdorff measure ${\mathcal{H}}^{1}$, such that ${\mathcal{S}t}\cup\mathcal{A}$ is connected. Such ${\mathcal{S}t}$ is proven to be acyclic. If ${\mathcal{S}t}$ itself is connected we call it Steiner tree for $\mathcal{A}$. The set $\mathcal{A}$ is called an input for the Steiner problem. See Section 2.3 for precise definition and properties of Steiner trees.
Following the paper [19], the question of finding non trivial examples of Steiner trees for infinitely many points was raised. Of course, it is easy to find solutions with infinitely many points by splitting a finite tree into infinitely many pieces. For example given any compact set of points on the real line, $A\subset\mathbb{R}$, the solution to the Steiner problem with input $A$ is clearly given by the set ${\mathcal{S}t}=[\min A,\max A]\setminus A$. If $A$ is compact the solution ${\mathcal{S}t}$ is composed of a countable number of disjoint open segments each with end points on $\partial A$. Much more difficult is to find an infinite Steiner tree which is indecomposable in the sense that it is not possible to split the tree ${\mathcal{S}t}$ into two connected pieces ${\mathcal{S}t}_{1}$ and ${\mathcal{S}t}_{2}$ so that both ${\mathcal{S}t}_{1}$ and ${\mathcal{S}t}_{2}$ are Steiner trees over a subset of the set $\mathcal{A}$ of terminal points.
Figure 1: The left part contains two Steiner trees connecting the vertices of a square; the right part illustrates an example of a self-similar solution $\Sigma(\lambda)$.
The first example of an infinite, indecomposable Steiner tree was given in [21]. However it was not self-similar and its input set $\mathcal{A}$ had zero Hausdorff dimension. The next natural question was to understand if the same result was true for the set of leaves of a self-similar binary tree $\Sigma(\lambda)$ composed of a segment of length $1$ which splits into two segments of length $\lambda$ each of which splits into two segments of length $\lambda^{2}$ and so on…The splitting points are regular tripods meaning that there are three segments joining in the point forming equal angles of size $32\pi/$ 3 (see Figure 1, right-hand side picture). The input set $\mathcal{A}(\lambda)$ is composed of a root point (the first end-point of the unit segment) and all the accumulation points of the smaller and smaller segments. Note that, if $\lambda<\frac{1}{2}$ the set $\mathcal{A}(\lambda)$ turns out to be a self-similar, fractal set, with Hausdorff dimension $-\frac{\ln 2}{\ln\lambda}>0$. Recently, in [6, 20], $\Sigma(\lambda)$ was proven to be the unique Steiner tree for the set $\mathcal{A}(\lambda)$ if $\lambda$ is small enough.
Clearly every connected subset $S$ of $\Sigma(\lambda)$ is the unique Steiner tree for the set of its endpoints and corner points (otherwise one can replace $S$ with a better competitor in $\Sigma(\lambda)$).
So $\Sigma(\lambda)$ can be regarded as a universal graph for Steiner trees in the sense that it contains a Steiner subtree with any given possible finite combinatorics. This fact was recently used in [1] to show that the set of $n$-point configurations in $\mathbb{R}^{d}$ with a unique Steiner tree is path-connected (as a subset of $\mathbb{R}^{dn}$).
In this paper we are interested in finding an explicit example of a Steiner tree connecting a countable number of points distributed along two lines forming an angle in the plane. The points will be equally distributed on the two sides of the angle in such a way that the distances of the points from the vertex of the angle are in a geometric progression.
A configuration like this was suggested in the Master’s Thesis of Letizia Ulivi, where an existence theorem for Steiner trees was proven in the case when the input set was considered to be compact and countable, see [22].
Examples of this kind are interesting because the set of terminals is distributed on the boundary of a convex set. In this setting the Steiner tree is splitting the convex set into many regions and hence determines a partition. The Steiner Problem becomes a problem of minimal partitions and can hence be studied in the framework of geometric measure theory by using calibrations (see [18, 16, 5]).
To our surprise we have found that, under suitable assumptions (Theorem 1), the Steiner tree for such a set of points is decomposable into rescaled copies of a Steiner tree connecting only 5 points. However by adding a pair of terminal points in a precise position, outside the geometric progression, we are able to force the Steiner tree to be indecomposable, see Remark 1.
The main difficulty in the proof is the fact that the set of locally minimal trees may have a very complicated structure (see Remark 3) and our methods heavily use the global optimality of a solution.
[…]
Figure 2: The input sets $\mathcal{A}_{0}=\{A_{\infty},A_{1},B_{1},A_{2},B_{2},\dots\}$, $\mathcal{A}_{0}=\mathcal{A}_{1}\cup[A_{0}B_{0}]$. Here $\alpha=\frac{\pi}{36}$ and $\lambda=\frac{1}{2}$.
Our setup.
Consider a positive $\lambda\leq\frac{1}{2}$ and an angle of size $2\alpha<\frac{\pi}{3}$ in the plane. Let $A_{\infty}$ be the vertex of the angle, and $A_{1},B_{1}$ be points on distinct sides of the angle at a unit distance from $A_{\infty}$ (so that the middle point of the segment $[A_{1}B_{1}]$ has distance $\cos\alpha$ from $A_{\infty}$). For $k\geq 1$, define $A_{k+1}\in[A_{k}A_{\infty}]$ and $B_{k+1}\in[B_{k}A_{\infty}]$ recursively by $|A_{k+1}A_{\infty}|=\lambda\cdot|A_{k}A_{\infty}|$ and $|B_{k+1}A_{\infty}|=\lambda\cdot|B_{k}A_{\infty}|$ (see Figure 2). Define $A_{0}$, $B_{0}$ on the rays $[A_{\infty}A_{1})$ and $[A_{\infty}B_{1})$, respectively, in such a way that
$|A_{0}A_{\infty}|=|B_{0}A_{\infty}|=\frac{1}{\lambda}-\frac{\tan(\alpha)}{\sqrt{3}\lambda}.$
Let
$\mathcal{A}_{1}=\mathcal{A}_{1}(\alpha,\lambda):=\{A_{\infty}\}\cup\bigcup_{n=1}^{\infty}\{A_{n},B_{n}\}$
and
$\mathcal{A}_{0}=\mathcal{A}_{0}(\alpha,\lambda):=\mathcal{A}_{1}(\alpha,\lambda)\cup[A_{0}B_{0}].$
The following theorem is the main result of the paper.
Theorem 1.
Suppose $0<\alpha<\frac{\pi}{6}$ and $0<\lambda\leq\frac{1}{2}$ are such that
$\sqrt{\lambda}<\frac{\cos\left(\frac{\pi}{3}+\alpha\right)}{\cos\left(\frac{\pi}{3}-\alpha\right)}.$ (1)
Then every solution ${\mathcal{S}t}_{1}$ to the Steiner problem for $\mathcal{A}_{1}$ is the union of full trees on 5 terminals (see Fig. 3). The length of a solution has the following explicit formula
${\mathcal{H}}^{1}({\mathcal{S}t}_{1})=\left|\cos\alpha+\sqrt{3}\sin\alpha+\frac{2\lambda}{1-\lambda^{2}}e^{\frac{\pi i}{6}}\sin\alpha+\frac{2\lambda^{2}}{1-\lambda^{2}}e^{-\frac{\pi i}{6}}\sin\alpha\right|.$
[…]
Figure 3: One of the solutions to the Steiner Problem for $\mathcal{A}_{1}$. See Theorem 1. Here $\alpha=\frac{\pi}{36}$ and $\lambda=\frac{1}{2}$.
Note that every Steiner tree for $\mathcal{A}_{1}$ is not full and so can be decomposed into solutions for 5-terminals subset of $\mathcal{A}_{1}$. Each of the solutions on 5 points can be mirrored with respect to the angle bisector, so we have $2^{\mathbb{N}}$ different solutions, all with the same length. Since every subtree of a Steiner tree is itself a Steiner tree with respect to its terminal points, we can state the following corollary, which holds in the framework of finite Steiner trees.
Corollary 1.
Suppose $\alpha$ and $\lambda$ are given with the assumptions of Theorem 1. Then the Steiner problem for the finite set $\{A_{1},\dots,A_{2k+1},B_{1},\dots,B_{2k}\}$ has $12^{k}-$ 1 different solutions each of which is the union of rescaled and mirrored copies of the Steiner tree on 5 terminals: $\{A_{1},A_{2},A_{3},B_{1},B_{2}\}$. The length also has an explicit formula.
We would like to slightly modify the input set $\mathcal{A}_{1}$ to obtain an indecomposable solution. It turns out that a simple way is to consider the input $\mathcal{A}_{0}$, which is obtained from $\mathcal{A}_{1}$ with the additional segment $[A_{0}B_{0}]$ (see definition above). We can then replace the segment with a single point $x$ placed in a precise position on the segment $[A_{0}B_{0}]$ or with two points $x_{1},x_{2}$ placed on the two sides of the angle in such a way that the Steiner tree has a triple point in the point $x$.
Theorem 2.
Suppose that $\alpha<\frac{\pi}{6}$ and $\lambda\leq 1/2$ satisfy (1). Then the following statements hold true.
  • (i)
    The Steiner problem for $\mathcal{A}_{0}$ has exactly 2 solutions ${\mathcal{S}t}_{0}^{1}$ and ${\mathcal{S}t}_{0}^{2}$ which are trees. The tree ${\mathcal{S}t}_{0}^{2}$ is the reflection of ${\mathcal{S}t}_{0}^{1}$ with respect to the angle bisector. Both solutions are indecomposable. The point $\{x\}={\mathcal{S}t}_{0}^{1}\cap[A_{0}B_{0}]$ (and the corresponding symmetric point ${\mathcal{S}t}_{0}^{2}\cap[A_{0}B_{0}]$) has a distance $\frac{1}{\lambda+1}\sin\alpha$ from the midpoint of $[A_{0}B_{0}]$.
  • (ii)
    The length of the solutions is
    ${\mathcal{H}}^{1}({\mathcal{S}t}_{0}^{1})={\mathcal{H}}^{1}({\mathcal{S}t}_{0}^{2})=\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}+\frac{\sqrt{3}}{1-\lambda}\sin\alpha.$
  • (iii)
    Each ${\mathcal{S}t}_{0}^{j}$, $j=1,2$, is a self-similar tree, in the sense that $f_{2}({\mathcal{S}t}_{0}^{j})\subset{\mathcal{S}t}_{0}^{j}$, where $f_{2}$ is the homothety with center $A_{\infty}$ and ratio $\lambda^{2}$.
[…]
Figure 4: One of the two Steiner trees ${\mathcal{S}t}_{0}^{1}$, ${\mathcal{S}t}_{0}^{2}$ for the set $\mathcal{A}_{0}$, see Theorem 2. The other solutions is the reflection with respect to the angle bisector. Here $\alpha=\frac{\pi}{36}$ and $\lambda=\frac{1}{2}$.
Remark 1.
If we consider a Steiner tree ${\mathcal{S}t}_{0}$ for $\mathcal{A}_{0}$ as stated in the previous theorem and insert a Steiner point in the vertex $x$ with two additional edges $[xA_{0}^{\prime}]$ and $[xB_{0}^{\prime}]$ so that $A_{0}^{\prime}$ and $B_{0}^{\prime}$ are on the edges of the angle $[A_{\infty},A_{1})$ and $[A_{\infty},B_{1})$, respectively, we obtain a full Steiner tree for the set of terminals $\mathcal{A}_{0}^{\prime}=\mathcal{A}_{1}\cup\{A_{0}^{\prime},B_{0}^{\prime}\}$ placed on the sides of the triangle $\triangle A_{0}^{\prime}A_{\infty}B_{0}^{\prime}$.
Structure of the paper.
Section 2 contains the notation and some general results we will use in the proofs. These results are well-known for a finite input and we translate them to an infinite setup. For the sake of completeness, the proofs are collected in the Appendix. In Section 3 we describe the rough structure of a Steiner tree for both input sets $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$. The proofs of Theorem 1 and 2 are given in Sections 4 and 5, respectively. In Section 6 we present a formulation to find a “locally minimal” competitor by means of a dynamical system. Section 7 collects open questions. In the Appendix we have collected the proofs of well known results.
2 Notation and basics
2.1 Notation
If $T$ is a subset of $\mathbb{R}^{d}$ we denote by $\overline{T}$, $\partial T$ and $\operatorname{conv}T$, respectively the closure, the boundary and the convex hull of $T$. For $\rho>0$ and $C\in\mathbb{R}^{d}$ we denote by $B_{\rho}(C)$ the open ball centered in $C$ with radius $\rho$. If $T$ is a set we denote by $B_{\rho}(T)=\bigcup_{C\in T}B_{\rho}(C)$ the open $\rho$-neighbourhood of $T$. We denote with ${\mathcal{H}}^{1}(T)$ the one-dimensional Hausdorff measure of the set $T$.
[…]
We let =$\mathbb{N}=\{1,2,3,\dots\}$ be the set of positive whole numbers.
2.2 Graphs and immersions
A graph $\Gamma$ is a pair $(V,E)$ where $V$ is any set and $E$ is a set such that every $e\in E$ is of the form $e=\{v,w\}$ with distinct $v,w\in V$. The elements of $V$ are called vertices of $\Gamma$ and the elements of $E$ are called edges. If $v$ is a vertex of $\Gamma$ we define the order (or degree) of $v$ as #$\#\{e\in E\colon v\in e\}$ i.e. the number of edges joining in $v$.
If each vertex of $\Gamma$ has finite order, we say that $\Gamma$ is a locally-finite graph. If $E$ and $V$ are both finite, we say that $\Gamma$ is a finite graph. A finite path $\gamma$ is an alternating sequence of vertices and edges: $v_{0},e_{1},v_{1},e_{2},v_{2},\dots,e_{n},v_{n}$ such that $e_{k}=\{v_{k-1},v_{k}\}$. If $v_{i}\neq v_{j}$ for all $(i,j)$ with $i<j$ and $(i,j)\neq(0,n)$ we say that the path is simple. A (simple) path with $v_{0}=v_{n}$ will be called a (simple) cycle. A path with no edges will be called trivial otherwise it will be called non-trivial. If for all pair of vertices $v,w\in V$ there is a finite path with first vertex $v$ and last vertex $w$ we say that the graph is connected. If there is no non-trivial cycle we say that the graph is a forest. A connected forest is called a tree.
Given $\varphi\colon V\to\mathbb{R}^{d}$ we say that the set
$S_{\varphi}=\bigcup_{\{v,w\}\in E}[\varphi(v)\,\varphi(w)]\subset\mathbb{R}^{d}$
is a geodesic immersion (or simply an immersion) of $\Gamma$ in $\mathbb{R}^{d}$. The points $\varphi(v)$ are called immersed vertices while the segments $[\varphi(v)\varphi(w)]$ with $\{v,w\}\in E$ are called immersed edges. The length of an immersed graph is defined by
$\ell(\varphi)=\sum_{\{v,w\}\in E}\lvert\varphi(v)\,\varphi(w)\rvert.$
Clearly one has $\ell(\varphi)\geq{\mathcal{H}}^{1}(S_{\varphi})$.
If the graph is finite, the map $\phi$ is injective and the segments of the immersion have pairwise disjoint interior (i.e. the intersection of $]\varphi(v)\,\varphi(w)[$ and $]\varphi(v^{\prime})\,\varphi(w^{\prime})[$ is empty whenever $\{v,w\}\neq\{v^{\prime},w^{\prime}\}$) we say that the immersion is an embedding.
In the following we will often refer to immersed graphs, immersed vertices and immersed edges simply as graphs, vertices and edges.
We will say that an immersed graph $S_{\varphi}$ is full* when the abstract graph is connected and all the convex angles between any two edges adjacent to the same vertex, are equal to $32\pi/$ 3. In particular in a full* tree all vertices have order at most three. The wind-rose of an immersed graph is the union of all lines passing through the origin and parallel to any edge of the graph. Clearly, in the planar case $d=2$, the wind-rose of a full* immersed graph comprises at most three different lines (so the corresponding directed segments have at most six directions).
We will say that an immersed graph $S_{\varphi}$ is full if it is full* and there are no vertices of order two.
2.3 General Steiner trees
If $S$ is a subset of a topological space, we say that $S$ is connected if it is not the disjoint union of two non-empty closed sets. We say that $S$ is path-connected if given any two points of $S$ there is a continuous curve joining them in $S$. We say that $S$ contains no loop if no subset of $S$ is homeomorphic to the circle $\mathbb{S}^{1}$. We say that $S$ is a topological tree if $S$ is connected and contains no loop.
Let $\mathcal{A}\subset\mathbb{R}^{d}$ and ${\mathcal{S}t}\subset\mathbb{R}^{d}$. We say that ${\mathcal{S}t}$ is a Steiner set for $\mathcal{A}$ if ${\mathcal{S}t}\cup\mathcal{A}$ is connected and ${\mathcal{H}}^{1}({\mathcal{S}t})\leq{\mathcal{H}}^{1}(S)$ whenever $S\subset\mathbb{R}^{d}$ and $S\cup\mathcal{A}$ is connected. We say that ${\mathcal{S}t}$ is a Steiner tree if ${\mathcal{S}t}$ is a Steiner set and is itself connected.
The following theorem collects some of the general results from [19].
Theorem 3.
If $\mathcal{A}\subset\mathbb{R}^{d}$ is a compact set then there exists a Steiner set ${\mathcal{S}t}$ for $\mathcal{A}$. Moreover if ${\mathcal{S}t}$ is a Steiner set for $\mathcal{A}$ and ${\mathcal{H}}^{1}({\mathcal{S}t})<+\infty$ then
  1. 1.
    ${\mathcal{S}t}\cup\mathcal{A}$ is compact;
  2. 2.
    ${\mathcal{S}t}\setminus\mathcal{A}$ has at most countable many connected components, and each of them has positive length;
  3. 3.
    $\overline{\mathcal{S}t}$ contains no loop (homeomorphic image of $\mathbb{S}^{1}$);
  4. 4.
    the closure of every connected component of ${\mathcal{S}t}$ is a topological tree with endpoints on $\mathcal{A}$ (in particular it has at most countable number of branching points) and has at most one endpoint on each connected component of $\mathcal{A}$;
  5. 5.
    if $\mathcal{A}$ is finite then $\overline{\mathcal{S}t}={\mathcal{S}t}\cup\mathcal{A}$ is an embedding of a finite tree;
  6. 6.
    for almost every $\varepsilon>0$ the set ${\mathcal{S}t}\setminus B_{\varepsilon}(\mathcal{A})$ is an embedding of a finite graph.
We need to strengthen the result of the previous theorem to the case when $\mathcal{A}$ is countable. In that case it is useful to prove that the Steiner set ${\mathcal{S}t}$ is locally finite also in the neighborhood of the isolated points of $\mathcal{A}$. The proof of the following statements can be found in the Appendix.
Lemma 1.
Let $x_{0}\in\mathcal{A}$ be an isolated point of the compact set $\mathcal{A}$. Let ${\mathcal{S}t}$ be a Steiner set for $\mathcal{A}$ with ${\mathcal{H}}^{1}({\mathcal{S}t})<+\infty$. Then there exists $\rho>0$ such that $\overline{{\mathcal{S}t}\cap B_{\rho}(x_{0})}$ is an embedding of a finite graph.
Corollary 2.
Let $\mathcal{A}\subset\mathbb{R}^{d}$ be a compact set. Let $\mathcal{A}^{\prime}$ be the set of accumulation points of $\mathcal{A}$ and $\mathcal{A}_{0}=\mathcal{A}\setminus\mathcal{A}^{\prime}$ be the set of isolated points of $\mathcal{A}$. Let ${\mathcal{S}t}$ be a Steiner set for $\mathcal{A}$ with ${\mathcal{H}}^{1}({\mathcal{S}t})<+\infty$. Then, for almost every $\rho>0$ the set ${\mathcal{S}t}\setminus B_{\rho}(\mathcal{A}^{\prime})$ is an embedding of a finite graph.
The previous corollary allows us to use the terminology of finite Steiner trees for a general Steiner tree ${\mathcal{S}t}$ with input $\mathcal{A}$. We will say that a point $x\in{\mathcal{S}t}\setminus\mathcal{A}$ is a Steiner point if there is $\rho>0$ such that ${\mathcal{S}t}\setminus B_{\rho}(\mathcal{A}^{\prime})$ is a finite embedded graph and $x$ is a vertex of this graph. We will say that a segment $[x,y]\subset{\mathcal{S}t}\setminus\mathcal{A}^{\prime}$ is an edge of the Steiner tree ${\mathcal{S}t}$ if there exists $\rho>0$ such that $[x,y]$ is an embedded edge of the finite embedded graph ${\mathcal{S}t}\setminus B_{\rho}(\mathcal{A}^{\prime})$.
Notice also that if ${\mathcal{S}t}$ is a Steiner set for $\mathcal{A}$ then ${\mathcal{S}t}\subset\operatorname{conv}\mathcal{A}$. In fact we know that the projection $\pi\colon\mathbb{R}^{d}\to\operatorname{conv}\mathcal{A}$ is a $1$-Lipschitz map, which implies that ${\mathcal{H}}^{1}(X)\leq{\mathcal{H}}^{1}(\pi(X))$ for every measurable set $X\subset\mathbb{R}^{d}$. and hence by letting ${\mathcal{S}t}^{\prime}=\pi({\mathcal{S}t})$ we have that ${\mathcal{S}t}^{\prime}\cup\mathcal{A}=\pi({\mathcal{S}t}\cup\mathcal{A})$ and hence ${\mathcal{S}t}^{\prime}\cup\mathcal{A}$ is connected and ${\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})\leq{\mathcal{H}}^{1}({\mathcal{S}t})$. This implies that ${\mathcal{S}t}^{\prime}$ is itself a Steiner set for $\mathcal{A}$. But then ${\mathcal{S}t}^{\prime}={\mathcal{S}t}$, otherwise one would have ${\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})<{\mathcal{H}}^{1}({\mathcal{S}t})$ which is a contradiction.
2.4 Maxwell’s length formula
In the following lemma we identify the Euclidean plane with the complex plane $\mathbb{C}$. The proof of the following Lemma can be found in the Appendix.
Lemma 2 (Maxwell-type formula).
Let $S_{\varphi}$ be a full*, finite, immersed tree and let $p_{1},p_{2},\dots,p_{n}\in\mathbb{C}$ be the vertices of order less than three. Suppose $n>1$. If $p_{k}$ is vertex of order one we define $c_{k}$ as the unit complex number representing the outer direction of the unique edge of $p_{k}$. If $p_{k}$ is a point of degree two then the two edges define an angle of $\frac{2}{3}\pi$ and we let $c_{k}$ be the unit complex number with the direction, away from $p_{k}$, of the third implied edge which would complete a regular tripod in $p_{k}$.
Then one can write the length of the tree as
$\ell(\phi)=\sum_{k=1}^{n}\bar{c_{k}}p_{k}.$ (2)
In particular the right hand side of (2) is a real number.
Remark 2.
If $\sum p_{k}$ converges absolutely then we can pass to the limit in (2). Clearly when $\{p_{1},p_{2},\dots\}=\mathcal{A}_{1}$ this sum converges absolutely if we choose $A_{\infty}$ to be the origin of the complex plane.
2.5 Melzak algorithm
The following lemma is an immediate corollary of Ptolemy’s theorem (see the Appendix for the proof).
Lemma 3 (Melzak’s reduction, [17]).
Let ${\mathcal{S}t}$ be a Steiner tree for $\mathcal{A}\subset\mathbb{R}^{2}$. If ${\mathcal{S}t}$ contains a branching point $q$ which is adjacent to two isolated points of $\mathcal{A}$, say $p_{1}$ and $p_{2}$, and the circle passing through $p_{1},p_{2}$ and $q$ contains no other point of ${\mathcal{S}t}$ apart from the two edges $[p_{1}q]$ and $[p_{2}q]$, then the length ${\mathcal{H}}^{1}({\mathcal{S}t})$ is equal to the length of the tree ${\mathcal{S}t}^{\prime}$ obtained from ${\mathcal{S}t}$ by removing the points $p_{1}$ and $p_{2}$ and the edges $[p_{1}q]$ and $[p_{2}q]$ and prolonging the remaining edge arriving in $q$ up to a new vertex $p$ such that $pp_{1}p_{2}$ is an equilateral triangle and (among the two possibilities) $p$ and $q$ are on opposite sides of the line $(p_{1},p_{2})$.
In the following we will call $p$ the Melzak point for $p_{1}$ and $p_{2}$.
2.6 A convexity argument
Let $G=(V,E)$ be an abstract graph and $\varphi$ an immersion of $G$ in $\mathbb{R}^{d}$ with $\ell(\varphi)<+\infty$. If $V_{0}\subset V$ is fixed, we say that the immersion $\varphi$ is locally minimal with $V_{0}$ fixed if there exists $\delta>0$ such that $\ell(\varphi)\leq\ell(\psi)$ for every map $\psi\colon V\to\mathbb{R}^{d}$ with $\psi(v)=\varphi(v)$ when $v\in V_{0}$ and $\lvert\psi(v)\,\psi(v)\rvert<\delta$ for all $v\in V$.
The following theorem shows that the directions of the edges of a locally minimal immersion are uniquely determined. Often this is enough to deduce the uniqueness of locally, and hence globally minimal immersion of a given graph.
Theorem 4.
Let $\Gamma=(V,E)$ be a graph and $\varphi,\psi\colon V\to\mathbb{R}^{d}$ be two locally minimal immersions of $G$ into $\mathbb{R}^{d}$ with fixed $V_{0}\subset V$. Suppose that $\ell(\varphi)<+\infty$. Then $\ell(\psi)=\ell(\varphi)$ and one has that the edges of $\varphi$ are parallel to the corresponding edges of $\psi$. More precisely, given any edge $\{v,w\}\in E$ either it has zero length in one of the immersions ($\varphi(v)=\varphi(w)$ or $\psi(v)=\psi(w)$) or there exists $t>0$ such that $\varphi(v)-\varphi(w)=t(\psi(v)-\psi(w))$.
3 Preliminary lemmas
Lemma 4.
Let $\mathcal{A}$ be a compact subset of the Euclidean plane and let $\mathcal{W}$ be a closed geometrical angle of size at least $\frac{2}{3}\pi$ such that $\mathcal{A}\cap\mathcal{W}=\emptyset$. Let ${\mathcal{S}t}$ be a Steiner tree for $\mathcal{A}$ with ${\mathcal{H}}^{1}({\mathcal{S}t})<\infty$. Then ${\mathcal{S}t}\cap\mathcal{W}$ is either a segment or an empty set.
Proof.
First of all we show that there are no Steiner points in $\mathcal{W}$. This is because in a Steiner point at least one of the three directions is contained in the angle $V$ and hence in that direction another Steiner point should be found (because there are no points of $\mathcal{A}$ in $V$), which is also inside the angle and further away from the vertex. These Steiner points should have an accumulation point on $A$ which is not possible since $A\cap\mathcal{W}=\emptyset$.
Since there are no Steiner points in $\mathcal{W}$ we can state that ${\mathcal{S}t}\cap\mathcal{W}$ is composed of line segments with end points on $\partial\mathcal{W}$. Assume that there are two such segments $[A_{1}B_{1}]$ and $[A_{2}B_{2}]$, where $A_{1},A_{2}$ belong to the same side of $\mathcal{W}$ while $B_{1},B_{2}$ are on the other side. Let $W$ be the vertex of the angle. Without loss of generality we assume that $|A_{1}B_{1}|\leq|A_{2}B_{2}|$. Then both $|A_{1}A_{2}|$ and $|B_{1}B_{2}|$ are shorter than $|A_{2}B_{2}|$, because the longest side of the triangle $A_{2}B_{2}W$ is $[A_{2}B_{2}]$ and $[A_{1}A_{2}],[B_{1}B_{2}]$ are subsets of $[A_{2}W],[B_{2}W]$, respectively.
Since ${\mathcal{S}t}$ is connected there is a path in ${\mathcal{S}t}\setminus\mathcal{W}$ which joins the two segments $[A_{1}B_{1}]$ and $[A_{2}B_{2}]$. Hence if we remove the segment $[A_{2}B_{2}]$ and replace it with $[A_{1}B_{1}]$ or $[A_{2}B_{2}]$ we are able to obtain a connected set which is shorter than the original tree ${\mathcal{S}t}$ but with the same terminals. This would contradict the minimality of ${\mathcal{S}t}$ and conclude the proof. ∎
Lemma 5.
Suppose that $\alpha<\frac{\pi}{6}$ and $\lambda$ satisfy (1). Let ${\mathcal{S}t}$ be a Steiner tree for the set $\mathcal{A}_{0}$ or $\mathcal{A}_{1}$. Then for all $k\in\mathbb{N}$ there exist two disjoint convex sets $\mathcal{C}_{k}$ and $\mathcal{D}_{k}$ such that $A_{j},B_{j}\in\mathcal{C}_{k}$ for all $j\leq k$, $A_{j},B_{j}\in\mathcal{D}_{k}$ for all $j>k$, and ${\mathcal{S}t}\setminus(\mathcal{C}_{k}\cup\mathcal{D}_{k})$ is a line segment with one end point on $\partial\mathcal{C}_{k}$ and one end point on $\partial\mathcal{D}_{k}$.
Proof.
Let us show that the assumptions of the lemma imply, for all $k\in\mathbb{N}$, the existence of points $W_{1}\in[B_{k}B_{k+1}]$ and $W_{2}\in[A_{k}A_{k+1}]$ such that the two angles $\angle A_{k}W_{1}A_{k+1}$ and $\angle B_{k}W_{2}B_{k+1}$ have size at least $32\pi/$ 3 (see Fig. 5). Let $W_{1}$ be the point in $[B_{k},B_{k+1}]$ such that $\triangle A_{k}A_{\infty}W_{1}$ is similar to $\triangle W_{1}A_{\infty}A_{k+1}$, which is equivalent to say that $|A_{\infty}A_{k}|\cdot|A_{\infty}A_{k+1}|=|A_{\infty}W_{1}|^{2}$. So $|A_{\infty}W_{1}|=\lambda^{k-1/2}$. Let $2\gamma:=\angle A_{k}W_{1}A_{k+1}$. By the sine rule for $\triangle A_{\infty}A_{k}W_{1}$
$\frac{\sqrt{\lambda}}{\sin\angle A_{\infty}A_{k}W_{1}}=\frac{1}{\sin\angle A_{\infty}W_{1}A_{k}}.$
Notice that since $\angle A_{\infty}A_{k}W_{1}=\angle A_{\infty}W_{1}A_{k+1}$ we have $\pi=2\alpha+2\gamma+2\angle A_{\infty}A_{k}W_{1}$ hence $\angle A_{\infty}A_{k}W_{1}=\frac{\pi}{2}-\alpha-\gamma$ and $\angle A_{\infty}W_{1}A_{k}=2\gamma+\angle A_{\infty}W_{1}A_{k+1}=2\gamma+\frac{\pi}{2}-\alpha-\gamma=\frac{\pi}{2}-\alpha+\gamma$. Thus
$\sqrt{\lambda}=\frac{\cos(\alpha+\gamma)}{\cos(\alpha-\gamma)}.$
The right-hand side is decreasing in $\gamma$ since
$\frac{d}{d\gamma}\frac{\cos(\alpha+\gamma)}{\cos(\alpha-\gamma)}=-\frac{\sin 2\alpha}{\cos^{2}(\alpha-\gamma)}.$
Hence (1) implies $\gamma>\frac{\pi}{3}$ and $\angle A_{k}W_{1}A_{k+1}>\frac{2\pi}{3}$. Define $W_{2}$ as the reflection of $W_{1}$ with respect to the angle bisector.
[…]
Figure 5: Definition of the point $W_{1}$ in Lemma 5
Now take $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ to be any angles of size $32\pi/$ 3 with vertices $W_{1}$ and $W_{2}$ contained in $\angle A_{k}W_{1}A_{k+1}$ and $\angle B_{k}W_{2}B_{k+1}$, respectively, so that $\mathcal{W}_{1},\mathcal{W}_{2}$ have empty intersection with the terminal set. Lemma 4 assures that each of the two angles $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ contains a single line segment of ${\mathcal{S}t}$: denote this segments respectively by $S_{1}$ and $S_{2}$. Since ${\mathcal{S}t}\subset\operatorname{conv}\mathcal{A}$ and the segment $[W_{1}W_{2}]$ splits $\operatorname{conv}\mathcal{A}$ into two parts, both segments $S_{1}$ and $S_{2}$ must intersect $[W_{1}W_{2}]$ and hence $S_{1}\cup S_{2}$ must be itself a single line segment. So the statement is proven by taking $\mathcal{C}_{k}$ and $\mathcal{D}_{k}$ to be the connected components of $\mathbb{R}^{2}\setminus(\mathcal{W}_{1}\cup\mathcal{W}_{2})$. ∎
Lemma 6.
Suppose that $\alpha<\frac{\pi}{6}$ and $\lambda\leq\frac{1}{2}$ satisfy (1).
  • (i)
    Any full* Steiner tree ${\mathcal{S}t}_{1}$ for $\mathcal{A}_{1}$ is uniquely determined by its wind-rose.
  • (ii)
    Any Steiner set for $\mathcal{A}_{0}$ is a tree.
  • (iii)
    Any full* Steiner tree ${\mathcal{S}t}_{0}$ for $\mathcal{A}_{0}$ is uniquely determined by ${\mathcal{S}t}_{0}\cap[A_{0}B_{0}]$, which is always a point.
  • (iv)
    Let ${\mathcal{S}t}={\mathcal{S}t}_{1}$ or ${\mathcal{S}t}={\mathcal{S}t}_{0}$. Then for every $k\in N$ there is a unique parallelogram $\mathcal{P}_{k}$ whose sides are parallel to the wind-rose of ${\mathcal{S}t}$.
  • (v)
    Every full finite component of a full* Steiner tree ${\mathcal{S}t}_{1}$ for $\mathcal{A}_{1}$ contains an odd number of terminals.
Proof.
Let us start with item (iv). In the case of the set ${\mathcal{S}t}={\mathcal{S}t}_{0}$ notice that the edge of ${\mathcal{S}t}_{0}$ touching the segment $[A_{0}B_{0}]$ must be parallel to the bisector and hence cannot be used to construct a parallelogram $\mathcal{P}_{k}$. The parallelogram is hence uniquely defined by the other two directions of the wind rose. In the case of ${\mathcal{S}t}={\mathcal{S}t}_{1}$ the proof is more involved. Let $\mathcal{C}_{1}$ be the convex set given by Lemma 5 applied to ${\mathcal{S}t}_{1}$ with $k=1$. The Lemma assures that $A_{1}$ and $B_{1}$ are the only terminals in $\mathcal{C}_{1}$ and that ${\mathcal{S}t}_{1}\cap\partial\mathcal{C}_{1}$ is a point. Since ${\mathcal{S}t}_{1}\subset\operatorname{conv}\mathcal{A}_{1}$ (see Section 2.3), points $A_{1}$ and $B_{1}$ have degree one in ${\mathcal{S}t}_{1}$, so ${\mathcal{S}t}_{1}\cap\mathcal{C}_{1}$ is a tripod. Thus $A_{1}$ and $B_{1}$ are adjacent to the same Steiner point, let us call it $S_{1}$. Consider the Melzak point $C$ for $A_{1}$ and $B_{1}$; clearly $C$ lies in the bisector. Put
$\mathcal{A}^{\prime}=\mathcal{A}_{1}\cup\{C\}\setminus\{A_{1},B_{1}\}.$
By Lemma 3 applied with $p_{1}=A_{1}$, $p_{2}=B_{1}$ and $q=C$ a full* tree ${\mathcal{S}t}_{1}$ corresponds to a full* tree ${\mathcal{S}t}^{\prime}$ for $\mathcal{A}^{\prime}$ with the same wind-rose.
Since ${\mathcal{S}t}^{\prime}\subset\operatorname{conv}\mathcal{A}^{\prime}$, the segment of ${\mathcal{S}t}^{\prime}$ incident to $C$ belongs to the angle $A_{2}CB_{2}$. Note that $\angle A_{2}CB_{2}<2\alpha=\angle A_{2}A_{\infty}B_{2}$ because $|A_{\infty}A_{2}|=|A_{\infty}B_{2}|<|A_{2}C|=|B_{2}C|$ since $\lambda\leq\frac{1}{2}$. Thus the wind-rose of ${\mathcal{S}t}^{\prime}$ contains a direction defining an angle smaller than $\alpha$ with the bisector. This means that either $A_{k}$ or $B_{k}$ has only two directions from the wind-rose which go inside $\operatorname{conv}\mathcal{A}$ and thus $\mathcal{P}_{k}$ is uniquely defined for ${\mathcal{S}t}_{1}$.
Denote the vertices of $\mathcal{P}_{k}$ by $A_{k}V_{k}B_{k}U_{k}$ for $k\geq 1$, where $V_{k}\in\operatorname{conv}\{A_{\infty},A_{k},B_{k}\}$. We claim that the parallelograms $\mathcal{P}_{k}$ are disjoint. Call $I$ the intersection of the segments $[A_{1}B_{2}]$ and $[A_{2}B_{1}]$. By symmetry $I$ belongs to the bisector. Then (1) implies $\angle A_{1}IA_{2}>\angle A_{1}W_{1}A_{2}\geq 2\pi/3$, where $W_{1}$ is defined in the proof of Lemma 5. So $\angle A_{1}IB_{1}=\angle A_{2}IB_{2}=\pi-\angle A_{1}IA_{2}<\pi/3$. Thus the domains $\{X\,|\,\angle A_{1}XB_{1}\geq 2\pi/3\}\supset\mathcal{P}_{1}$ and $\{Y\,|\,\angle A_{2}YB_{2}\geq 2\pi/3\}\supset\mathcal{P}_{2}$ are separated by the perpendicular to the bisector at $I$ and hence $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are disjoint. By similarity this implies that all parallelograms $\mathcal{P}_{k}$ are disjoint.
Now we are going to prove that there is a single edge in ${\mathcal{S}t}_{0}$ which has a point on the segment $[A_{0}B_{0}]$ which implies item (ii). Assume the contrary, then ${\mathcal{S}t}_{0}\setminus[A_{0}B_{0}]$ has at least 2 components. One of them contains $A_{\infty}$, so its length is at least
$\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}.$
Also, the length of every component cannot be smaller than the distance between $[A_{0}B_{0}]$ and $\mathcal{A}_{1}$ which is
$\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}-\cos\alpha.$
Summing up
${\mathcal{H}}^{1}({\mathcal{S}t}_{0})\geq 2\frac{\cos\alpha}{\lambda}-2\frac{\sin\alpha}{\sqrt{3}\lambda}-\cos\alpha.$
Under the assumption (1) this is greater than the length of the example we have [^1] Formally we have a logical loop here, but the calculation of length in a given example is independent with any logical arguments., whose length is
$\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}+\frac{\sqrt{3}}{1-\lambda}\sin\alpha;$
this is equivalent to
$\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}-\cos\alpha\geq\frac{\sqrt{3}}{1-\lambda}\sin\alpha$ (3)
which is shown in the end of the Appendix.
Let us prove items (i) and (iii). Recall that $S_{1}$ is the Steiner point adjacent to $A_{1}$ and $B_{1}$ in ${\mathcal{S}t}_{1}$, and also $S_{1}$ is a vertex of $\mathcal{P}_{1}$ (both for ${\mathcal{S}t}_{0}$ and ${\mathcal{S}t}_{1}$).
Now consider, in ${\mathcal{S}t}_{j}$, the simple path $P_{k}$ from $A_{k}$ to $B_{k}$ for $k>1$ (see Fig. 6). The number of connected components in ${\mathcal{S}t}_{j}\setminus P_{k}$ is equal to the number of branching points in $P_{k}$ plus the number of terminals of degree 2 in $\{A_{k},B_{k}\}$. By Lemma 5 the path $P_{k}$ contains exactly two points from $\overline{{\mathcal{S}t}_{j}\setminus P_{k}}$ (such points are either branching or terminal). We call these points $S_{k}$, $T_{k}$, in a way that $S_{k}\in[A_{k}V_{k}]\cup[V_{k}B_{k}]$ and $T_{k}\in[A_{k}U_{k}]\cup[U_{k}B_{k}]$. The corresponding edges have opposite directions in ${\mathcal{S}t}_{j}\setminus P_{k}$. Thus $P_{k}\subset\mathcal{P}_{k}$.
Reasoning by induction on $k$ the point $S_{k}$ determines $T_{k+1}$ and $T_{k+1}$ determines $S_{k+1}$. This proves items (i) and (iii).
[…]
Figure 6: Behaviour of ${\mathcal{S}t}_{1}$ and ${\mathcal{S}t}_{0}$ around the terminals $A_{k}$ and $B_{k}$.
Now we proceed with item (v). Note that in the construction we cannot have $\{S_{k},T_{k}\}=\{A_{k},B_{k}\}$, because otherwise we would have $P_{k}=[A_{k}B_{k}]$, but this edge has a direction which is not represented in the wind-rose. Thus the full component ${\mathcal{S}t}_{1}^{\prime}$ of ${\mathcal{S}t}_{1}$ containing $A_{1}$ has an odd (or an infinite) number of terminals. Note that the terminal set of $\overline{{\mathcal{S}t}_{1}\setminus{\mathcal{S}t}_{1}^{\prime}}$ is similar to $\mathcal{A}$ with a scale factor $\lambda^{k}$, so a reduction implies item (iii).
4 Proof of Theorem 1
Lemma 7.
If $\alpha<\frac{\pi}{6}$, and $\lambda$ satisfy (1), then there exists a full* Steiner tree for $\mathcal{A}_{1}$.
Proof.
Let ${\mathcal{S}t}_{1}$ be a solution to the Steiner problem for $\mathcal{A}_{1}$. Let ${\mathcal{S}t}^{\prime}$ be the full component of ${\mathcal{S}t}_{1}$ containing $A_{1}$. If ${\mathcal{S}t}^{\prime}$ is infinite then it contains every segment between $\mathcal{C}_{k}$ and $\mathcal{D}_{k}$, defined in Lemma 5, so every other full component should be a segment $[A_{k}B_{k}]$ which is impossible. This means ${\mathcal{S}t}^{\prime}={\mathcal{S}t}_{1}$ and we are done.
If ${\mathcal{S}t}^{\prime}$ is finite, then by Lemma 6(v) ${\mathcal{S}t}^{\prime}$ has an odd number of terminals and without loss of generality the largest index of a terminal of ${\mathcal{S}t}^{\prime}$ is reached on $A_{k+1}$. Then $\mathcal{A}_{k+1}:=\mathcal{A}\setminus\{A_{1},\dots,A_{k}\}\setminus\{B_{1},\dots,B_{k}\}$ is similar to $\mathcal{A}_{1}$, i.e.
$f_{k}(\mathcal{A}_{1})=\mathcal{A}_{k+1},$
where $f_{k}$ is the homothety with center $A_{\infty}$ and ratio $\lambda^{k}$. So
${\mathcal{H}}^{1}({\mathcal{S}t}_{k+1})=\lambda^{k}{\mathcal{H}}^{1}({\mathcal{S}t}_{1}),$
where ${\mathcal{S}t}_{k+1}$ is a Steiner tree for $\mathcal{A}_{k+1}$ and
${\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})=(1-\lambda^{k}){\mathcal{H}}^{1}({\mathcal{S}t}_{1}).$
Let ${\mathcal{S}t}^{*}$ be the union of $f_{k}^{j}({\mathcal{S}t}^{\prime})$ over non-negative integers $j$. Note that ${\mathcal{S}t}^{*}$ connects $\mathcal{A}_{1}$ and
${\mathcal{H}}^{1}({\mathcal{S}t}^{*})=\sum_{j=0}^{\infty}\lambda^{kj}{\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})=\frac{1}{1-\lambda^{k}}{\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})={\mathcal{H}}^{1}({\mathcal{S}t}_{1}).$
Thus ${\mathcal{S}t}^{*}$ is a solution to the Steiner problem for $\mathcal{A}_{1}$. By construction it is full*, so the lemma is proved. ∎
Lemma 8.
The length of a full* Steiner tree for $\mathcal{A}_{1}$ is equal to
$e^{-\beta i}\left(\cos\alpha+\sqrt{3}\sin\alpha+ae^{\frac{\pi i}{6}}+be^{-\frac{\pi i}{6}}\right)=\left|\cos\alpha+\sqrt{3}\sin\alpha+ae^{\frac{\pi i}{6}}+be^{-\frac{\pi i}{6}}\right|,$
where $a=2\sin\alpha\sum_{j\in J}\lambda^{j}$, $b=2\sin\alpha\sum_{j\in\mathbb{N}\setminus J}\lambda^{j}$ for some $J\subset\mathbb{N}$.
Proof.
Consider a complex coordinate system, determined by $A_{\infty}=0$, $A_{1}=\cos\alpha+i\sin\alpha$, $B_{1}=\cos\alpha-i\sin\alpha$. Let $C$ be a Melzak point of $A_{1}$ and $B_{1}$; clearly, $C=\cos\alpha+\sqrt{3}\sin\alpha$ and
$A_{k}=\lambda^{k-1}(\cos\alpha+i\sin\alpha),\quad\quad B_{k}=\lambda^{k-1}(\cos\alpha-i\sin\alpha).$
Denote by $\beta$ the angle smallest angle between an edge of the tree and the real axis; by the argument in the proof of Lemma 6(iv), one has $\beta\in[0,\alpha]$. Now we apply Lemma 2. By Melzak algorithm we can replace $A_{1}$ and $B_{1}$ with $C$. The unitary coefficient at $C$ is $e^{-\beta i}$, at $A_{k}$ is $e^{(-2\pi/3-\beta)i}$ or $e^{(-\pi/3-\beta)i}$ and at $B_{k}$ is $e^{(\pi/3-\beta)i}$ or $e^{(2\pi/3-\beta)i}$, respectively. By the proof of Lemma 6(i) the sum of the coefficients at $A_{k}$ and $B_{k}$ is zero for every $k>1$. So the summands with indices $k$ are either equal to
$2\sin\alpha\,e^{(-2\pi/3-\beta)i}\lambda^{k-1}i=2\sin\alpha\,e^{(-\pi/6-\beta)i}\lambda^{k-1}$
or to
$2\sin\alpha\,e^{(-\pi/3-\beta)i}\lambda^{k-1}i=2\sin\alpha\,e^{(\pi/6-\beta)i}\lambda^{k-1}.$
Proper naming of $a$ and $b$ gives us the desired result. ∎
Lemma 9.
Suppose that $\lambda\leq\frac{1}{2}$ and $\alpha<\frac{\pi}{6}$ satisfy (1). If ${\mathcal{H}}^{1}({\mathcal{S}t}_{1}^{1})={\mathcal{H}}^{1}({\mathcal{S}t}_{1}^{2})$ for distinct full* Steiner trees ${\mathcal{S}t}_{1}^{1}$ and ${\mathcal{S}t}_{1}^{2}$ for $\mathcal{A}_{1}$, then ${\mathcal{S}t}_{1}^{2}$ is the reflection of ${\mathcal{S}t}_{1}^{2}$ with respect to the angle bisector.
Proof.
Assume the contrary. Lemma 8 gives
$\left|\cos\alpha+\sqrt{3}\sin\alpha+a_{1}e^{\frac{\pi i}{6}}+b_{1}e^{-\frac{\pi i}{6}}\right|=\left|\cos\alpha+\sqrt{3}\sin\alpha+a_{2}e^{\frac{\pi i}{6}}+b_{2}e^{-\frac{\pi i}{6}}\right|.$
Put $c=\cos\alpha+\sqrt{3}\sin\alpha$ and $L_{j}=\cos\alpha+\sqrt{3}\sin\alpha+a_{j}e^{\frac{\pi i}{6}}+b_{j}e^{-\frac{\pi i}{6}}$, $j=1,2$. Then
$L_{j}\cdot\bar{L_{j}}=c^{2}+c(a_{j}+b_{j})(e^{\frac{\pi i}{6}}+e^{-\frac{\pi i}{6}})+a_{j}^{2}+b_{j}^{2}+a_{j}b_{j}(e^{\frac{\pi i}{3}}+e^{-\frac{\pi i}{3}})=c^{2}+\sqrt{3}c(a_{j}+b_{j})+(a_{j}+b_{j})^{2}-a_{j}b_{j}.$ (4)
Recall that $a_{j}+b_{j}=2\sin\alpha\cdot\frac{\lambda}{1-\lambda}$, so $L_{1}\cdot\bar{L_{1}}=L_{2}\cdot\bar{L_{2}}$ implies $a_{1}b_{1}=a_{2}b_{2}$, hence we have either $a_{1}=a_{2}$ and $b_{1}=b_{2}$ or $a_{1}=b_{2}$ and $a_{2}=b_{1}$. Without loss of generality we may assume that we are in the first case.
Let $a_{1}$ and $a_{2}$ be the sums of $\lambda^{k}$ over $k\in J_{1}$ and $k\in J_{2}$, respectively. Since $\lambda\leq\frac{1}{2}$ the sets of indices $J_{1}$ and $J_{2}$ coincide, say $J_{1}=J_{2}:=J$. Define an abstract graph $G=(V,E)$ with vertices
$V:=\{A_{k},B_{k},S_{k},T_{k+1}\,|\,k\in\mathbb{N}\}$
and edges
$E:=\{(A_{1},S_{1}),(S_{1},B_{1})\}\cup\{(S_{k},T_{k+1})\,|\,k\in\mathbb{N}\}\cup$
$\{(A_{k+1},S_{k+1}),(B_{k+1},T_{k+1})\,|\,k\in J\}\cup\{(A_{k+1},T_{k+1}),(B_{k+1},S_{k+1})\,|\,k\in\mathbb{N}\setminus J\}.$
Recall that ${\mathcal{S}t}_{1}^{1}$ and ${\mathcal{S}t}_{1}^{2}$ are Steiner trees, so they are global (and hence local) minima of $\ell_{G}(\cdot)$. By Theorem 4 for $G$ and $V_{0}=\mathcal{A}_{1}\setminus\{A_{\infty}\}$ the trees ${\mathcal{S}t}_{1}^{1}$ and ${\mathcal{S}t}_{1}^{2}$ have a common wind-rose and by Lemma 6(i) they coincide. Thus in the case $a_{1}=b_{2}$ and $a_{2}=b_{1}$ the trees ${\mathcal{S}t}_{1}^{1}$ and ${\mathcal{S}t}_{1}^{2}$ are symmetric. ∎
Now we are ready to prove Theorem 1.
Remark 3.
Here we have shown that a full* Steiner tree for $\mathcal{A}_{1}$ is determined (up to the reflection) by knowing which of the two possible directions of the wind-rose is used in each vertex $A_{k}$ (this determines the set $J\subset\mathbb{N}$). We don’t know which are the sets $J$ that correspond to a locally minimal tree: the answer may depend on $\alpha$ and $\lambda$ in some complicated way (we discuss it in Section 6). Fortunately, a relatively simple argument allows us to find a very strict condition on $J$ which must hold for an absolute minimimer.
[…]
Figure 7: The construction used in the proof of Theorem 1.
Consider the complex coordinate system from Lemma 8.
Let ${\mathcal{S}t}_{1}$ be a full* Steiner tree for $\mathcal{A}_{1}$ and let $\mathcal{P}_{k}$ be the parallelogram $A_{k}V_{k}B_{k}U_{k}$ defined in Lemma 6(iv). Without loss of generality suppose that $S_{1}=V_{1}$ lies in the lower half-plane (otherwise consider the reflection of ${\mathcal{S}t}_{1}$). Thus every $U_{j}$ lies in the upper half-plane and every $V_{j}$ belongs to the lower half-plane. Clearly for every $j$ the segments $[U_{j}A_{j}]$ lie in the upper half-plane and $[V_{j}B_{j}]$ lie in the lower half-plane. We need the following observations.
  1. 1.
    One has $|V_{j}B_{j}|=|U_{j}A_{j}|<|U_{j}B_{j}|=|A_{j}V_{j}|$ for every $j$ as $U_{j}$ belongs to the upper half-plane and the inequality $|XA_{j}|<|XB_{j}|$ holds for every point $X$ from the upper half-plane.
  2. 2.
    By Lemma 3 the point $T_{2}$ lies in the lower half-plane as $T_{2}$ belongs to the ray $[C,S_{1})$, where $C$ is the Melzak point of $A_{1}$ and $B_{1}$ (recall that $C$ belongs to the real axis) and thus $S_{1}$ belongs to the lower half-plane. Then we have $T_{2}\in[V_{2}B_{2}]$.
  3. 3.
    If $T_{j}$ belongs to the upper half-plane then $T_{j+1}$ belongs to the lower half-plane as the ray $[S_{j}T_{j+1})$ is co-directed with the ray $[S_{1}T_{2})\subset[CT_{2})$ and $S_{j}$ belongs to $[V_{j}B_{j}]$ and hence to the lower half-plane.
Lemma 8 says that
${\mathcal{H}}^{1}({\mathcal{S}t}_{1})=e^{-\beta i}\left(\cos\alpha+\sqrt{3}\sin\alpha+ae^{\frac{\pi i}{6}}+be^{-\frac{\pi i}{6}}\right)=\left|\cos\alpha+\sqrt{3}\sin\alpha+ae^{\frac{\pi i}{6}}+be^{-\frac{\pi i}{6}}\right|,$
where $a=2\sin\alpha\sum_{j\in J}\lambda^{j}$, $b=2\sin\alpha\sum_{j\in\mathbb{N}\setminus J}\lambda^{j}$ for some $J\subset\mathbb{N}$. Looking at equation (4) one notices that since $a+b$ is fixed (independent of $J$) the total length is smaller when $ab$ is larger, or, which is the same, when $|a-b|$ is smaller. By observation 2, one has $1\in J$. Observation 3 says that $j$ and $j+1$ cannot belong to $\mathbb{N}\setminus J$ simultaneously. So $|a-b|$ is minimal for $J=\{2k-1\,|\,k\in\mathbb{N}\}$. The construction described in the proof of Lemma 7 started with a full tree for $A_{1},A_{2},A_{3},B_{1},B_{2}$ shows that such $J$ is admissible. So the following lemma concludes the proof.
Lemma 10.
Suppose that $\lambda\leq\frac{1}{2}$ and $\alpha<\pi/6$ satisfy the condition (1). Then every Steiner tree for $\mathcal{A}_{5}:=\{A_{1},A_{2},A_{3},B_{1},B_{2}\}$ is full.
Proof.
Assume the contrary, then a Steiner tree ${\mathcal{S}t}_{5}$ contains a terminal $X$ of degree 2. Recall that the Steiner tree is contained in $\operatorname{conv}\mathcal{A}_{5}=A_{3}A_{1}B_{1}B_{2}$ (see Figure 3). Clearly, $\angle A_{3}A_{1}B_{1}=\angle A_{1}B_{1}B_{2}=\pi/2-\alpha<2\pi/3$ and $\angle B_{2}A_{3}A_{1}<\angle B_{3}A_{3}A_{1}=\pi/2+\alpha<2\pi/3$. Thus $X$ is either $A_{2}$ or $B_{2}$. Repeating the argument of Lemma 5 we conclude that $A_{1}$ and $B_{1}$ are adjacent to the same Steiner point $Y$. Now if both $A_{2}$ and $B_{2}$ have degree 2, then ${\mathcal{S}t}_{5}$ contains segment $[A_{2}B_{2}]$, which is impossible since $\angle A_{3}A_{2}B_{2},\angle A_{3}B_{2}A_{2}\leq\pi/2-\alpha$.
Now suppose that ${\mathcal{S}t}_{5}$ contains a full component which is an edge $[XZ]$. Clearly $[XZ]$ cannot split $\operatorname{conv}\mathcal{A}_{5}$ because $[XZ]\neq[A_{2}B_{2}]$, so $Z=A_{3}$. Since $|A_{2}A_{3}|<|B_{2}A_{3}|$, in this case ${\mathcal{S}t}_{5}$ is a union of $[A_{2}A_{3}]$ and a symmetric solution for $\{A_{1},B_{1},A_{2},B_{2}\}$. But then $\angle A_{3}A_{2}W=2\pi/3-\alpha$, where $W$ is the second Steiner point, a contradiction.
In the remaining case ${\mathcal{S}t}_{5}$ is a union of two similar tripods, which is longer by (4). ∎
5 Proof of Theorem 2
Proof of item (i)..
Clearly $\mathcal{A}_{0}$ is a compact set, so the Steiner problem has a solution. By Lemma 6(ii) every Steiner set for $\mathcal{A}_{0}$ is a Steiner tree. Suppose that ${\mathcal{S}t}_{0}$ is a Steiner tree for $\mathcal{A}_{0}$; obviously ${\mathcal{H}}^{1}({\mathcal{S}t}_{0})<\infty$. Denote the point $\overline{{\mathcal{S}t}_{0}}\cap[A_{0}B_{0}]$ by $x$.
Lemma 11.
${\mathcal{S}t}_{0}$ is a full* tree.
Proof.
Consider the full* subtree of ${\mathcal{S}t}_{0}$ containing $x$ which is maximal with respect to inclusion. Call it ${\mathcal{S}t}_{0}^{\prime}$. Since the segment of ${\mathcal{S}t}_{0}$ containing $x$ is orthogonal to $[A_{0}B_{0}]$, the wind rose of ${\mathcal{S}t}_{0}^{\prime}$ contains the direction of the bisector of the angle. Suppose, by contradiction that ${\mathcal{S}t}_{0}\setminus{\mathcal{S}t}_{0}^{\prime}$ is non-empty. Then ${\mathcal{S}t}_{0}^{\prime}$ contains a terminal (without loss of generality suppose it is $A_{k}$) which has degree 2 in ${\mathcal{S}t}_{0}$ and degree 1 in ${\mathcal{S}t}_{0}^{\prime}$.
Let $[A_{k}y]$, $[A_{k}z]$ be the edges incident to $A_{k}$, where $z$ is a vertex of ${\mathcal{S}t}_{0}^{\prime}$ and $y$ is a vertex of ${\mathcal{S}t}_{0}$ but not of ${\mathcal{S}t}_{0}^{\prime}$. Since ${\mathcal{S}t}_{0}^{\prime}$ is maximal $\angle yA_{k}z>2\pi/3$ and hence the segment $[A_{k}z]$ must be parallel to the bisector.
If $y$ is a terminal point then by Lemma 5 $y\in\{A_{k+1},B_{k+1}\}$. By Lemma 5 the edge $[A_{k}y]$ contains the segment of ${\mathcal{S}t}_{0}$ connecting $\mathcal{C}_{k}$ and $\mathcal{D}_{k}$. Hence $A_{k}$ and $B_{k}$ belong to the same component of ${\mathcal{S}t}_{0}\setminus[yA_{k}[$. By Lemma 5 the tree ${\mathcal{S}t}_{0}$ contains unique segment connecting $\mathcal{C}_{k-1}$ and $\mathcal{D}_{k-1}$, so the path connecting $A_{k}$ and $B_{k}$ in ${\mathcal{S}t}_{0}$ has exactly one Steiner point, which should be $z$. This is impossible since $[A_{k}z]$ is parallel to the bisector.
Thus $y$ is a branching point and so $[yB_{k}]\subset{\mathcal{S}t}_{0}$; recall that the angle $yA_{k}z$ is strictly greater than $32\pi/$ 3. Let ${\mathcal{S}t}$ be the union of ${\mathcal{S}t}_{0}\setminus{\mathcal{S}t}_{0}^{\prime}$ and the reflection of ${\mathcal{S}t}_{0}^{\prime}$. Let $z^{\prime}$ be the reflection of $z$, then $B_{k}z^{\prime}$ is also parallel to the bisector. Note that ${\mathcal{H}}^{1}({\mathcal{S}t})={\mathcal{H}}^{1}({\mathcal{S}t}_{0})$, so ${\mathcal{S}t}$ is also a solution to the Steiner problem for $\mathcal{A}$. But
$\angle yB_{k}z^{\prime}=2\pi-\angle zA_{k}y-\angle A_{k}yB_{k}<2\pi/3,$
so ${\mathcal{S}t}$ is not a Steiner tree, contradiction.
Lemma 12.
${\mathcal{S}t}_{0}$ is a full tree.
Proof.
By Lemma 11 ${\mathcal{S}t}_{0}$ is a full* tree. Let ${\mathcal{S}t}_{0}^{\prime}$ be the full component containing $x$. Consider a terminal of ${\mathcal{S}t}_{0}^{\prime}$ with degree $2$ (without loss of generality it is $A_{k}$). Then the structure of the other full component ${\mathcal{S}t}$, containing $A_{k}$ is uniquely determined $S_{k}$ is the vertex of the rhombus $\mathcal{R}_{k}$. Thus ${\mathcal{S}t}$ is a Steiner tree for 4 terminals $A_{k},B_{k},A_{k+1},B_{k+1}$ (and its branching points are $S_{k},T_{k+1}$). But
$\angle A_{\infty}A_{k+1}T_{k+1}=\angle A_{\infty}B_{k+1}T_{k+1}=2\pi/3-\alpha$
which means that both $A_{k+1}$ and $B_{k+1}$ cannot be a point of degree 2 in ${\mathcal{S}t}_{0}$. This is a contradiction. ∎
Lemma 13.
Consecutive horizontal parts of ${\mathcal{S}t}_{0}$ lie on the opposite sides of the bisector of the angle.
Proof.
We repeat the proof of Lemma 6(iii). Every $\mathcal{P}_{i}$ is a rhombus with two vertices at the bisector of the angle. By Lemma 12 no segment of ${\mathcal{S}t}_{0}$ belongs to the bisector. Thus $S_{i+1}$ and $T_{i+1}$ always belong to the opposite sides of the bisector. Since $S_{i}T_{i+1}$ is parallel to the bisector for every $i$, we are done. ∎
Lemma 14.
Consider a complex coordinate system, such that $A_{\infty}=0$, $A_{1}=\cos\alpha+i\sin\alpha$, $B_{1}=\cos\alpha-i\sin\alpha$. Let $x$ be the point where $\overline{{\mathcal{S}t}_{0}}$ touches $[A_{0}B_{0}]$ and suppose that $\lvert A_{0}x\rvert\leq\lvert B_{0}x\rvert$. Then
${\mathcal{H}}^{1}({\mathcal{S}t}_{0})=x+\left(\frac{\sqrt{3}}{1-\lambda}-\frac{i}{1+\lambda}\right)\sin\alpha$
where $x$ is the complex number representing the point $x$. In the case when $\lvert A_{0}x\rvert\leq\lvert B_{0}x\rvert$ we obtain the same formula with $x$ replaced by $\bar{x}$.
Proof.
We are going to use the length formula (2):
${\mathcal{H}}^{1}({\mathcal{S}t}_{0})=\sum_{j=1}^{\infty}\bar{c_{j}}A_{j}+\sum_{j=1}^{\infty}\bar{d_{j}}B_{j}+1\cdot x.$ (5)
By Lemma 12 every terminal has degree 1. If we suppose that $\lvert A_{0}x\rvert\leq\lvert B_{0}x\rvert$ the first horizontal edge of ${\mathcal{S}t}_{0}$ starting from $x$ is going to hit the upper part of the first rhombus $\mathcal{R}_{1}$. This means that the first Steiner point $S_{1}$ is connected to the upper terminal $A_{1}$ and has a direction $e^{2\pi i/3}$. The edge connected to the terminal $B_{1}$ has the opposite direction. By Lemma 13 the next horizontal edge is below the angle bisector and the direction of the edges on the terminals $A_{2}$ and $B_{2}$ are rotated by $-\pi/3$. The angles keep alternating between even and odd vertices, so, one has the following coefficients $c_{k},d_{k}$ in (5): $c_{2k-1}=e^{2\pi i/3}$, $c_{2k}=e^{\pi i/3}$, $d_{2k-1}=e^{-\pi i/3}$ and $d_{2k}=e^{-2\pi i/3}$, $k\in\mathbb{N}$. The coefficient for $x$ is $1$. Recall that $A_{j}=\lambda^{j}e^{i\alpha}$, $B_{j}=\lambda^{j}e^{-i\alpha}$. For odd $j$ one has $\overline{c_{j}}A_{j}+\overline{d_{j}}B_{j}=2\lambda^{j}\sin\alpha\cdot i\cdot e^{-2\pi i/3}$ while for even $j$ one obtains $2\lambda^{j}\sin\alpha\cdot i\cdot e^{-\pi i/3}$. Thus (5) becomes:
${\mathcal{H}}^{1}({\mathcal{S}t}_{0})$ $=x+2(1+\lambda^{2}+\lambda^{4}\dots)ie^{-2\pi i/3}\sin\alpha+2\lambda(1+\lambda^{2}+\lambda^{4}+\dots)ie^{-\pi i/3}\sin\alpha$
$=x+\left(\frac{2e^{-2\pi i/3}}{1-\lambda^{2}}+\frac{2\lambda e^{-\pi i/3}}{1-\lambda^{2}}\right)i\sin\alpha=x+\left(\frac{2e^{-\pi i/6}}{1-\lambda^{2}}+\frac{2\lambda e^{\pi i/6}}{1-\lambda^{2}}\right)\sin\alpha$
$=x+\left(\frac{\sqrt{3}}{1-\lambda}-\frac{i}{1+\lambda}\right)\sin\alpha.$
Now we are ready to finish the proof of the theorem. By Lemma 12 ${\mathcal{S}t}_{0}$ is full*. Lemma 14 gives a formula for the length, which should be a real number. Thus
$\Im x=\pm\frac{1}{1+\lambda}\sin\alpha.$
This determines the point $x$ proving item (i). Recall that $\Re x=\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}$, so
${\mathcal{H}}^{1}({\mathcal{S}t}_{0})=\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}+\frac{\sqrt{3}}{1-\lambda}\sin\alpha$
and item (ii) is also proven.
We are left with the proof of item (iii). Clearly ${\mathcal{S}t}_{0}^{\prime}={\mathcal{S}t}_{0}\cap\operatorname{conv}f_{2}(\mathcal{A}_{0})$ connects $f_{2}(\mathcal{A}_{0})$. Note that ${\mathcal{S}t}_{0}^{\prime}$ is a full tree so we may apply Lemma 2 to compute ${\mathcal{H}}^{1}({\mathcal{S}t}_{0}^{\prime})$. This computation is essentially contained in the computation in the proof of Lemma 14; the only difference is that we omit two starting terms in geometric progressions. Thus we obtain
${\mathcal{H}}^{1}\left({\mathcal{S}t}^{\prime}_{0}\right)=\lambda^{2}{\mathcal{H}}^{1}({\mathcal{S}t}_{0}).$
It means that ${\mathcal{S}t}_{0}\cap f_{2}([A_{0}B_{0}])$ is a Steiner tree for $f_{2}(\mathcal{A}_{0})$ and Lemma 13 gives $f_{2}(x)={\mathcal{S}t}_{0}\cap f_{2}([A_{0}B_{0}])$. Also $f_{-1}(\mathcal{P}_{1})\cap[A_{0}B_{0}]=f_{-1}(V_{1})$ so $f_{-1}(\mathcal{P}_{1})$ does not intersect ${\mathcal{S}t}_{0}$. Thus $f_{2}({\mathcal{S}t}_{0})$ does not intersect $\mathcal{P}_{2}$ and so $f_{2}({\mathcal{S}t}_{0})\subset{\mathcal{S}t}_{0}$. ∎
6 The dynamical system
Lemma 6(i) assures that any full* tree with terminals on the set $\mathcal{A}_{1}$ is uniquely determined by its wind-rose. If one knows the position of the line containing the long edge joining the first two parallelograms $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ then the whole tree is determined. Since the points of $\mathcal{A}_{1}$ are placed in geometric progression, each subsequent step of the tree construction is subject to the same recurrence equation with the appropriate rescaling.
It turns out that the resulting discrete dynamical system is conjugate to the inverse of a 2-interval piecewise contraction, whose law can be written as
$F_{\lambda,\delta}:x\in[0,1]\to\{\lambda x+\delta\},$
[…]
Suppose ${\mathcal{S}t}$ is a full* tree satisfying the assumptions of Lemma 5. We know that the segments of ${\mathcal{S}t}$ outside the parallelograms $\mathcal{P}_{i}$ are all parallel to each other. Let $e_{1}$ be the unit vector parallel to these segments pointing from $T_{2}$ towards $S_{1}$ in Figure 7. Let $e_{2}$ and $e_{3}$ be the two unit vectors completing the wind-rose of ${\mathcal{S}t}$ so that $e_{1}$, $e_{2}$, $e_{3}$ have zero sum and are counter-clockwise oriented ($e_{2}$ is pointing from $T_{2}$ towards $S_{2}$ and $e_{3}$ is pointing from $T_{2}$ towards $B_{2}$). It is convenient to use a hexagonal (barycentric) coordinate system centered in $A_{\infty}$. A point $p$ is an equivalent class of triples $(u,v,w)$ under the relation $(u,v,w)=(u+t,v+t,w+t)$ which corresponds to the representation $p=ue_{1}+ve_{2}+we_{3}$. Usually we will choose the representant with $v+w=0$.
Using the notation of the proof Lemma 6 let $(l,\delta,-\delta)$ be the hexagonal coordinates of the point $C_{1}$ which is the center of the parallelogram $\mathcal{P}_{1}$. In general we have $\delta\neq 0$ because the direction $e_{1}$ is not parallel to the angle bisector. Let $2a:=|A_{1}V_{1}|=|U_{1}B_{1}|$ and $2b:=|A_{1}U_{1}|=|V_{1}B_{1}|$. One has:
$A_{1}$ $=C_{1}+(0,b,-a)=(l,b+\delta,-a-\delta)=\left(l+\frac{a-b}{2},\delta+\frac{a+b}{2},-\delta-\frac{a+b}{2}\right),$
$B_{1}$ $=C_{1}+(0,-b,a)=(l,\delta-b,a-\delta)=\left(l-\frac{a-b}{2},\delta-\frac{a+b}{2},\frac{a+b}{2}-\delta\right),$
$U_{1}$ $=C_{1}+(0,-b,-a)=(l,\delta-b,-a-\delta)=\left(l+\frac{a+b}{2},\delta+\frac{a-b}{2},-\delta-\frac{a-b}{2}\right),$
$V_{1}$ $=C_{1}+(0,b,a)=(l,\delta+b,a-\delta)=\left(l-\frac{a+b}{2},\delta-\frac{a-b}{2},\frac{a-b}{2}-\delta\right).$
Then define $C_{k}=\lambda^{k-1}C_{1}$, $A_{k}=\lambda^{k-1}A_{1}$, $B_{k}=\lambda^{k-1}B_{1}$, $U_{k}=\lambda^{k-1}U_{1}$ and $V_{k}=\lambda^{k-1}V_{1}$.
The line segments $[S_{k-1}T_{k}]$ connecting the parallelograms $\mathcal{P}_{k-1}$ and $\mathcal{P}_{k}$ identified by Lemma 5 are parallel to the direction $e_{1}$ hence have coordinates of the form $(\cdot,\mu_{k-1},-\mu_{k-1})$. We are going to find a recurrence relation between $\mu_{k-1}$ and $\mu_{k}$. Lemma 6(i) gives a rule to obtain $\mu_{k}$ from $\mu_{k-1}$, or to show that there is no corresponding $\mu_{k}$. Since the prolongation of the segment between the parallelograms $\mathcal{P}_{k-1}$ and $\mathcal{P}_{k}$ must pass between the points $A_{k}$ and $B_{k}$, one should have
$\left(\delta-\frac{a+b}{2}\right)\lambda^{k-1}\leq\mu_{k-1}\leq\left(\delta+\frac{a+b}{2}\right)\lambda^{k-1}.$
We consider a renormalized version of $\mu_{k}$ by defining
$\nu_{k}:=\frac{1}{2}+\frac{\lambda^{-k}\mu_{k}-\delta}{a+b},\qquad\mbox{i.e.}\qquad\mu_{k}=\lambda^{k}\left((a+b)\left(\nu_{k}-\frac{1}{2}\right)+\delta\right)$ (6)
so that the previous conditions on $\mu_{k}$ become $\nu_{k}\in[0,1]$.
Suppose, without loss of generality, that $b>a$. Define
$q^{\pm}$ $:=\frac{1}{2}+\frac{(1-\lambda)\delta-a}{\lambda(a+b)}\pm\frac{1}{2\lambda}$
$t_{1}$ $:=\lambda(1-q^{+}),\quad t^{*}:=\frac{a}{a+b},\quad t_{2}:=-\lambda q^{-}.$
Lemma 15.
One has $\nu_{k+1}=g(\nu_{k})$ with
$g(t)=\begin{cases}\frac{t}{\lambda}+q^{+}&\text{if }0\leq t<t^{*},\\ \frac{t}{\lambda}+q^{-}&\text{if }t^{*}<t\leq 1.\end{cases}$ (7)
Moreover $g([0,t_{1}]\cup[t_{2},1])=[0,1]$, $g(0)=g(1)$, $g$ injective on $\left]0,1\right[\setminus\{t^{*}\}$. If we restrict the values of $g$ to the interval $[0,1]$, the inverse function, $g^{-1}\colon[0,1]\setminus\{q^{+}\}\to[0,1]$ is uniquely defined and one has
$g^{-1}(t)=\left\{\lambda t+t_{2}\right\}$
where $\left\{x\right\}=x-\lfloor x\rfloor$ is the fractional part of $x$.
[…]
Figure 8: Possible behaviour of a locally minimal tree in the parallelogram $\mathcal{P}_{k}$.
Proof.
We have only two possible behaviours at $\mathcal{P}_{k}$ (they are drawn in blue and green in Fig. 8). When the input line coming from the right hand side hits the segment $[A_{k},U_{k}]$ (green case) the line goes down, on the opposite side $[B_{k},V_{k}]$. The output line is obtained adding the vector $2a\lambda^{k}e_{3}$ to the input. In hexagonal coordinates:
$(t,\mu_{k},-\mu_{k})\mapsto(t,\mu_{k},2a\lambda^{k}-\mu_{k})=(t-a\lambda^{k},\mu_{k}-a\lambda^{k},a\lambda^{k}-\mu_{k}).$
In this case we obtain $\mu_{k+1}=\mu_{k}-a\lambda^{k}$. This happens when the line $(t,\mu_{k},-\mu_{k})$ passes above $U_{k}=\lambda^{k}U_{1}$, that is, when $\mu_{k}>\lambda^{k}\left(\delta+\frac{a-b}{2}\right)$.
Otherwise, if $\mu_{k}<\lambda^{k}\left(\delta+\frac{a-b}{2}\right)$, the input line hits the segment $[B_{k},U_{k}]$ and it goes up, towards the opposite side $[A_{k},V_{k}]$. In this case we add the vector $2b\lambda^{k}e_{2}$ obtaining:
$(t,\mu_{k},-\mu_{k})\mapsto(t,\mu_{k}+2b\lambda^{k},-\mu_{k})=(t-b\lambda^{k},\mu_{k}+b\lambda^{k},-\mu_{k}-b\lambda^{k}).$
Hence $\mu_{k+1}=\mu_{k}+b\lambda^{k}$.
A line hitting the extremal points $A_{k}$ or $B_{k}$ will cause the tree to have a terminal of degree two (a finite full subtree). A line hitting the point $U_{k}$ would split hitting both terminals $A_{k}$ and $B_{k}$: by the argument of Lemma 6(v) this is not possible in a minimal Steiner tree.
In conclusion, we have $\lambda^{-1-k}\mu_{k+1}=h(\lambda^{-k}\mu_{k})$ with
$h(x)=\begin{cases}\frac{x+b}{\lambda}&\text{if $-\frac{a+b}{2}<x-\delta<\frac{a-b}{2}$}\\ \\ \frac{x-a}{\lambda}&\text{if $\frac{a-b}{2}<x-\delta<\frac{a+b}{2}$}.\end{cases}$
To find the renormalized recurrence relation we put $\nu_{k}=\psi(\lambda^{-k}\mu_{k})$ with
$\psi(s)=\frac{1}{2}+\frac{s-\delta}{a+b},\qquad\psi^{-1}(t)=(2t-1)\frac{a+b}{2}+\delta$
whence $\nu_{k+1}=g(\nu_{k})$ with $g=\psi\circ h\circ\psi^{-1}$. With a straightforward computation one finds:
$g(t)=\psi(h(\psi^{-1}(t)))=\frac{t}{\lambda}+\frac{1}{2}+\frac{(1-\lambda)\delta-a}{\lambda(a+b)}\pm\frac{1}{2\lambda}\qquad\text{if $t\lessgtr\frac{a}{a+b}$}.$
This gives (7).
One has $g(t)=\frac{t}{\lambda}+q^{\pm}$. Hence the inverse function has the form $g^{-1}(s)=\lambda(s-q^{\pm})$. For $q^{+}<s\leq 1$ one has $g^{-1}(s)=\lambda(s-q^{+})$. Since $10<\lambda<$ 1 it is clear that in this case $10<\lambda(s-q^{+})<$ 1 hence $g^{-1}(s)=\{\lambda(s-q^{+})\}$ for $q^{+}<s\leq 1$.
If $0\leq s<q^{+}$ one has $g^{-1}(s)=\lambda(s-q^{-})$. But notice that $q^{-}=q^{+}-\frac{1}{\lambda}$ hence $\lambda(s-q^{-})=\lambda(s-q^{+})+1$. Notice that in this case $-1<\lambda(s-qk)<0$ and hence, also in this case, $g^{-1}(s)=\lambda(s-q^{+})+1=\{\lambda(s-q^{+})\}$.
Now consider the dynamical system $([0,1],g)$ which depends only on $\beta$. First, recall our proof scheme for $\mathcal{A}=\mathcal{A}_{1}$ (Theorem 1). We restrict the attention on full* candidates, then we show that the length is determined by $J$ and the simplest cut of cases gave us that an optimal $J$ is $2\mathbb{N}$.
Suppose now that one wants to check an existence of a competitor for a given $J$. Then $\beta$ can be easily determined as the argument in Lemma 8. Lemma 15 explicitly defines the dynamics and it remains to check if the trajectory started in $\nu_{1}$ is infinite. For a full* solution from Theorem 1 the corresponding trajectory is 2-periodic with the period $\nu_{1},0$ or $\nu_{1},1$. Note that this gives a way to finish the proof of Theorem 1 avoiding Lemma 10.
Also, let us note that the trajectory which corresponds to a locally minimal full tree for $\mathcal{A}_{1}$ is not periodic.
One can force a full* tree to have a given $\beta$ by addition a segment $[A_{\beta}B_{\beta}]$ with a proper slope to $\mathcal{A}_{1}$. In the particular case of $\beta=0$ we add the segment $[A_{0}B_{0}]$ to obtain the input $\mathcal{A}_{0}$; other parameters are $a=b$, $\delta=0$ and $g^{-1}(t)=\{\lambda t+1-\lambda/2\}$. Steiner trees ${\mathcal{S}t}_{0}^{1}$ and ${\mathcal{S}t}_{0}^{2}$ from Theorem 2 again correspond to the 2-periodic points of $([0,1],g)$, namely $t=\frac{\lambda}{2(\lambda+1)}$ and $t=\frac{\lambda+2}{2(\lambda+1)}$. However for a general $\beta$ there is no simple analogue of Lemma 13, which again makes the situation more intricate.
7 Open questions
Gilbert–Pollak conjecture on the Steiner ratio.
A minimum spanning tree is a shortest connection of a given set $A$ of points in the plane by segments with endpoints in $A$. It is well known that a minimum spanning tree can be found in a polynomial time.
The Steiner ratio is the infimum of the ratio of the total length of Steiner tree to the minimum spanning tree for a finite set of points in the Euclidean plane. The most known open problem on Steiner trees is whether the planar Steiner ratio is reached on an equilateral triangle (and so is equal to $\frac{\sqrt{3}}{2}$). This question is known as Gilbert–Pollak conjecture; despite the fact that some “proofs” are published this conjecture is open, see [13]. Graham [12] offered 1000\$ for the positive answer.
Analogues results in space.
Steiner trees in a Euclidean $d$-dimensional space have the same local structure as in the plane, id est every branching point has degree 3 and moreover three adjacent segments have pairwise angles $32\pi/$ 3 and are coplanar. However a global structure is much more complicated. In particular a full (and full*) tree may have no wind-rose. Up to our knowledge all known series of explicit solutions are planar.
In particular, we are interested in a solution to the Steiner problem for an input which is a mix of a regular $n$-gon with the setup of this paper. Consider a regular $3$-dimensional $n$-gon in the plane $x=1$ centered at $(1,0,0)$; let $\mathcal{Q}$ be the set of its vertices. Then the input is the union of all $f^{k}(\mathcal{Q})$, $k\in\mathbb{N}$, where $f$ is a homothety with the center in the origin and a small enough scale factor $\lambda$.
Some of our tools (in particular, Lemma 4) can be translated to the $d$-dimensional setup.
The number of solutions.
It is known that the total number of full Steiner topologies with $n$ terminals is
$\frac{(2n-4)!}{2^{n-2}(n-2)!}.$
How many of them may simultaneously give a shortest solution to the Steiner problem? In all known examples (for instance, ladders or Corollary 1) the number is at most exponential in $n$ while the number of topologies has a factorial order of growth. Note that the answer may heavily depend on the dimension.
On the other hand, for $n\geq 4$ the set of planar $n$-point configurations with multiple solution to Steiner problem has Hausdorff dimension $12n-$ 1 (as a subset of $(\mathbb{R}^{2})^{n}$), see [1]. It is interesting to find the corresponding dimension in a $d$-dimensional Euclidean space.
Density questions.
The following questions make sense both in two dimensions and in higher dimensions. Some of them seem not very difficult and can be considered as exercises. Let ${\mathcal{S}t}$ be a Steiner tree and suppose $B_{r}(x)$ (or another good set) contain no terminal point of ${\mathcal{S}t}$. What is the largest possible number of branching points of ${\mathcal{S}t}$ in $B_{r}(x)$? What is the maximum of ${\mathcal{H}}^{1}({\mathcal{S}t}\cap B_{r}(x))/r$?
Notice that in this paper we proved that in the case of a Steiner tree with terminals on the boundary of a triangle we can have infinitely many Steiner points in the interior.
Let $\mathcal{A}$ be a finite subset of $\partial B_{1}(0)$ (or of another good curve). What is the largest possible $\sharp({\mathcal{S}t}\cap\partial B_{1/2}(0))$? What is the largest possible number of branching points of ${\mathcal{S}t}$ lying inside $B_{1/2}(0)$? What is the upper bound for ${\mathcal{H}}^{1}({\mathcal{S}t}\cap B_{1/2}(0))$? Similar questions can be asked for any radius in place of $\frac{1}{2}$.
Acknowledgments.
We thank Pavel Prozorov for finding some gaps in the initial proof. We are grateful to Alexey Gordeev for some computer simulations and to Carlo Carminati for bringing the 2-interval piecewise contraction to our attention. Yana Teplitskaya is supported by the Simons Foundation grant 601941, GD.
Appendix: proofs of auxiliary and well-known results
Proof of Lemma 1.
If $\mathcal{A}=\{x_{0}\}$ the result is trivial. Without loss of generality, we can hence suppose that $\mathcal{A}$ contains at least two points.
Claim 1: there exists $\varepsilon>0$ such that every connected component of ${\mathcal{S}t}\setminus\mathcal{A}$ which intersect $\overline{B_{\varepsilon}(x_{0})}$ contains $x_{0}$. Let $\rho>0$ be such that $\overline{B_{2\rho}(x_{0})}\cap\mathcal{A}=\{x_{0}\}$. If a connected component $S$ of ${\mathcal{S}t}\setminus\mathcal{A}$ touches $\overline{B_{\rho}(x_{0})}$ but does not contain $x_{0}$, then it must touch both $\partial B_{\rho}(x_{0})$ and $\partial B_{2\rho}(x_{0})$. Being $S$ connected we would have ${\mathcal{H}}^{1}(S)\geq\rho$. Since ${\mathcal{H}}^{1}({\mathcal{S}t})<+\infty$, we conclude that only a finite number of connected components of ${\mathcal{S}t}\setminus\mathcal{A}$ can possibly intersect $\overline{B_{\rho}(x_{0})}$ without containing $x_{0}$. Take $\varepsilon$ smaller than the minimum distance of $x_{0}$ from each of these components to get the claim.
Claim 2: let $S_{0}$ be the connected component of $\overline{{\mathcal{S}t}}$ which contains $x_{0}$. Then $\overline{{\mathcal{S}t}\cap B_{\varepsilon}(x_{0})}=S_{0}\cap\overline{B_{\varepsilon}(x_{0})}$. Suppose by contradiction that there exists $x\in\overline{{\mathcal{S}t}\cap B_{\varepsilon}(x_{0})}\setminus S_{0}$. Let $S$ be the connected component of $\overline{{\mathcal{S}t}}$ which contains $x$ and let $S^{\prime}$ be the connected component of ${\mathcal{S}t}\setminus\mathcal{A}$ which contains $x$. Clearly $S^{\prime}\subset S$ and $x\in S^{\prime}\cap\overline{B_{\varepsilon}(x_{0})}$. This is in contradiction with Claim 1.
Claim 3: $S_{0}\cap\partial B_{t}(x_{0})\neq\emptyset$ for all $t<\varepsilon$. Suppose by contradiction that $\partial B_{t}(x_{0})\cap S_{0}=\emptyset$ for some $t<\varepsilon$. Then $S_{0}\subset B_{t}(x_{0})$ because $x_{0}\in S_{0}$ and $S_{0}$ is connected. This means that $S_{0}$ is also a connected component of ${\mathcal{S}t}\cup\mathcal{A}$ and this is not possible because ${\mathcal{S}t}\cup\mathcal{A}$ is connected by assumption and $\mathcal{A}$ is assumed to have some other point outside $B_{\varepsilon}(x_{0})$.
By Theorem 3, there exists $\rho<\varepsilon$ such that the set $\Sigma_{\rho}=\overline{{\mathcal{S}t}\setminus B_{\rho}(\mathcal{A})}$ is an embedding of a finite graph. Clearly $\Sigma_{\rho}\cap S_{0}\neq\emptyset$ because $S_{0}\subset\overline{{\mathcal{S}t}}$ touches $\partial B_{\rho}(x_{0})$, $\Sigma_{\rho}\supset{\mathcal{S}t}\cap\partial B_{\rho}(x_{0})$, $\overline{{\mathcal{S}t}}\setminus{\mathcal{S}t}\subset\mathcal{A}$ (Theorem 3) and $\partial B_{\rho}\cap\mathcal{A}=\emptyset$.
Hence $\Sigma=\Sigma_{0}\cup S_{0}$ is connected and contains $x_{0}$. Also $\Sigma\cap\partial B_{\rho}(x_{0})=\Sigma_{0}\cap\partial B_{\rho}(x_{0})$ is finite (by Theorem 3). The set $S=\Sigma\cap\overline{B_{\rho}}$ is compact and has a finite number of connected components because each connected component has touches $\Sigma\cap\partial B_{\rho}(x_{0})$, which is finite. Let $C$ be a connected component of $S$. We claim that $C$ is a Steiner set for $C\cap\partial B_{\rho}$, otherwise a Steiner set $C^{\prime}$ for $C\cap\partial B_{\rho}$ could be used to construct a competitor ${\mathcal{S}t}^{\prime}={\mathcal{S}t}\setminus C\cup C^{\prime}$ and obtain the contradiction ${\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})<{\mathcal{H}}^{1}({\mathcal{S}t})$. Since $C\cap\partial B_{\rho}\subset\Sigma\cap\partial B_{\rho}$ is finite, the set $S$ is a finite embedded tree (Theorem 3) and hence $S$ is an embedding of a finite graph. ∎
Proof of Corollary 2.
For almost every $\rho>0$ Theorem 3 assures that ${\mathcal{S}t}\setminus B_{\rho}(\mathcal{A})$ is an embedding of a finite graph. The set $\mathcal{A}\setminus B_{\rho}(\mathcal{A}^{\prime})$ is finite because it contains only isolated points without any accumulation point. For each $x_{0}$ in such finite set we can apply the previous lemma to find a neighborhood of $x_{0}$ where ${\mathcal{S}t}$ is an embedding of a finite graph. If $\varepsilon$ smaller than the minimum of the radii of these neighborhoods, we can apply, again, Theorem 3 to find that also ${\mathcal{S}t}\setminus B_{\varepsilon}(\mathcal{A})$ is an embedding of a finite graph and hence deduce that ${\mathcal{S}t}\setminus B_{\rho}(\mathcal{A}^{\prime})$ is an embedding of a finite graph. ∎
Proof of Lemma 2.
Let $v_{1},\ldots,v_{n+t}\in\mathbb{C}$ be the vertices of ${\mathcal{S}t}$ in the complex plane (they include the terminal points $v_{k}=p_{k}$ for $k=1,\dots,n$ and $t\leq n-2$ branching points) and let $E$ be the set of edges of ${\mathcal{S}t}$. Let $u(z):=\frac{z}{\lvert z\rvert}=e^{i\arg z}$ be the unit number representing the direction of the complex number $z$ so that $\lvert z\rvert=z\cdot\overline{u(z)}$. One has
${\mathcal{H}}^{1}({\mathcal{S}t})$ $=\sum_{(k,j)\in E}|v_{k}-v_{j}|=\sum_{(k,j)\in E}(v_{k}-v_{j})\cdot\overline{u(v_{k}-v_{j})}=\sum_{(k,j)\in E}\left(v_{k}\cdot\overline{u(v_{k}-v_{j})}+v_{j}\cdot\overline{u(v_{j}-v_{k})}\right)$
$=\sum_{k=1}^{n+t}v_{k}\cdot\sum_{(k,j)\in E}\overline{u(v_{k}-v_{j})}.$
If $v_{k}$ is a branching point then the corresponding sum of directions is zero, so we have only a sum over the terminal points. ∎
Proof of Lemma 3.
Clearly, $\angle p_{1}pp_{2}=\pi/3$, $\angle p_{1}qp_{2}=2\pi/3$ and the line $(p_{1},p_{2})$ separates $p$ and $q$. Hence $p,p_{1},p_{2}$ and $q$ are cocircular. Then $\angle p_{1}qp=\angle p_{1}p_{2}p=\pi/3$, so $[pq]$ is a continuation of the remaining edge adjacent to $q$ in ${\mathcal{S}t}$. By the Ptolemy’s theorem
$|pq|\cdot|p_{1}p_{2}|=|p_{1}q|\cdot|pp_{2}|+|p_{2}q|\cdot|pp_{1}|,$
so $|pq|=|p_{1}q|+|p_{2}q|$ and ${\mathcal{H}}^{1}({\mathcal{S}t}^{\prime})={\mathcal{H}}^{1}({\mathcal{S}t})$. ∎
Proof of Theorem 4.
We know that the function $x\mapsto\lvert x\rvert$ is a convex function defined on $\mathbb{R}^{d}$. Hence it turns out that the functional $\ell$ is convex on the vector space $(\mathbb{R}^{d})^{V}$ of all immersions $\varphi\colon V\to\mathbb{R}^{d}$. So, the real function $F(t)=\ell(t\varphi+(1-t)\psi)$ is convex for $t\in[0,1]$ and if $\varphi$ is a local minimum for $\ell$ it turns out that $t=0$ is a local minimum for $F$. Also, if $v\in V_{0}$, then $\varphi(v)=\psi(v)=(t\varphi+(1-t)\psi)\ (v)$, so $\ell$ is defined at $t\varphi+(1-t)\psi$. This easily implies that $F$ is constant and hence
$t\ell(\varphi)+(1-t)\ell(\psi)=\ell(t\varphi+(1-t)\psi)$ (8)
for every $t\in[0,1]$. Thus $\ell(\varphi)=\ell(\psi)$.
But since the inequality (8) is true for subgraph of $G$, in particular for every summand in the formula, hence equality in(8) holds in each such summand, hence for all $v$ and $w$ which are joined by an edge of $E$ one has
$t\lvert\varphi(v)-\varphi(w)\rvert+(1-t)\lvert\psi(v)-\psi(w)\rvert=\lvert t(\varphi(v)-\varphi(w))+(1-t)(\psi(v)-\psi(w))\rvert\qquad\forall t\in[0,1].$
But we notice that the equality $t\lvert x\rvert+(1-t)\lvert x\rvert=\lvert tx+(1-t)y\rvert$ holds for all $t\in[0,1]$ only when $x$ is a positive multiple of $y$ or when $x=0$ or $y=0$ since the function $\lvert x\rvert$ is linear only on the half-lines through the origin.
Letting $x=\varphi(v)-\varphi(w)$ and $y=\psi(v)-\psi(w)$ we obtain the claim of the theorem. ∎
Proof of (3).
We claimed that
$\frac{\cos\alpha}{\lambda}-\frac{\sin\alpha}{\sqrt{3}\lambda}-\cos\alpha\geq\frac{\sqrt{3}}{1-\lambda}\sin\alpha.$
This is equivalent to
$\cos\alpha\left(\frac{1}{\lambda}-1\right)>\sin\alpha\left(\frac{1}{\sqrt{3}\lambda}+\frac{\sqrt{3}}{1-\lambda}\right)$
and
$\cot\alpha>\frac{1+2\lambda}{\sqrt{3}(1-\lambda)^{2}}.$
One has
$\frac{d}{d\lambda}\frac{1+2\lambda}{\sqrt{3}(1-\lambda)^{2}}=\frac{2(2+\lambda)}{\sqrt{3}(1-\lambda)^{3}}$
which is positive for $20<\lambda\leq 1/$ 2. Cotangent is decreasing, so in view of (1) it is enough to examine
$\lambda=\left(\frac{\cos\left(\frac{\pi}{3}+\alpha\right)}{\cos\left(\frac{\pi}{3}-\alpha\right)}\right)^{2}=\left(\frac{\cot\alpha-\sqrt{3}}{\cot\alpha+\sqrt{3}}\right)^{2}.$
Put $t=\cot\alpha$, then the substitution gives that it is enough to show that
$\frac{-\sqrt{3}t^{4}+44t^{3}-2\sqrt{3}t^{2}-12t-9\sqrt{3}}{48t^{2}}>0$
for $t\in[\sqrt{3},3\sqrt{3}+2\sqrt{6}]$, were the ends correspond to $\alpha=\pi/6$ and $\lambda=1/2$, respectively. One can check that $-\sqrt{3}t^{4}+44t^{3}-2\sqrt{3}t^{2}-12t-9\sqrt{3}$ is increasing on this interval and so the minimal value is $96\sqrt{3}$ at $t=\sqrt{3}$.